What is the next term of the following sequence? 12, 6, 3, ...
hear r the answers
A) 1 1/2
B) 3/4
C) 0

Answer by checkley79(3341)

X/2 IS THE SOLUTION.
12/2=6
6/2=3
3/2=1.5
12, 6, 3, 1.5